9927 gallons of product is calculated to have been recovered according to one TSD facility where we have been sending the product/water mixture recovered from the storm drains. We have not received results from the second TSD facility.
Storm Drain Acitivities:
All sweeps have been recovered between manholes #2 and #3 near the pond. Divers were unable to fit down the manhole but the contractors were able to remove the last snagged sweeep with a grappling hook mounted on a jet rod. The sweep came out with some diesel staining, indicating very little product remains.
On site activities:
A hole in the planter next to the above ground storage tanks was created with a "super sucker". What appears to be pure diesel product began to flow into the hole. They began this operation on Friday and have been withdrawing product all day.
EPA START contractor continues to serve as overall safety officer.
Our operational period runs from 0900 Friday to 0900 Monday.
Work will continue over the weekend sucking up product from the hole in the planter. Work will also consist of exploratory soil removal to the extent of contamination.
EPA has placed some sorbent pads in three up-stream locations in the storm drains to remain over the weekend. Pads will be removed, bagged, and head space will be monitored for VOCs to compare as background with that of pads and sweeps removed downstream. A sampling plan will be prepared over the weekend and sent to Kathy Parker for review.
|